OpenStudy (anonymous):

Approximate the real number solutions to the polynomial function f(x) = x3 - 2x2 - 16x + 32 .

OpenStudy (anonymous):

f(x)=x^3-2x^2-16x+32 =x^2(x-2)-16(x-2) =(x-2)(x^2-16) =(x-2){(X)^2-(4)^2} =(x-2)(x-4)(x+4) f(x)=0 gives either x-2=0,x=2 or x-4=0,x=4 or x+4=0,x=-4 solution is x=2,4,-4